(2) A boat with a maximum speed with a width d. The river is flowing with speed V. (a) Traveling in a straight line (i.e perpendicular to the shore) how long does it take the boat to et to a point directly across? (b) What is the fastest possible crossing time to any point on the opposite shore? [20 pts] v (relative to the water) is on one shore of a river
Answer #1

(a)

Time taken by the boat to reach to a point directly across,

Time = width of river / speed of boat

t = d / v

(b)

For fastest possible crossing time,

speed of boat, v' = sqrt (v2 + V2)

t = d / v'

t = d /  sqrt (v2 + V2)